### v1

# A high-order recombination algorithm for weak approximation of stochastic differential equations

Abstract.
The authors present an algorithm for the application of the high-order recombination method first introduced by Terry Lyons and Christian Litterer in “high-order recombination and an application to cubature on Wiener space” (Ann. Appl. Probab 22(4):1301-1327, 2012), to real-world problems in mathematical finance. They also give numerical examples showing that the algorithm presented successfully avoids the explosive increase in the number of support of the measure that achieves high-order approximations.
Key words and phrases:
computational finance, stochastic differential equation, high-order discretisation method, cubature on Wiener space, weak approximation, tree-based method, Heston model
1991 Mathematics Subject Classification:
91G60, 60H35, 65C20, 65C30, 65C05, 68U20, 62P05
1. Introduction
In [16, 17] they show that the probability measure occurring in weak approximations of a stochastic differential equation (SDE) can be made to avoid an exponential increase in the number of points of its support with respect to accuracy by using a method of transforming the probability measure of a given finite probability space into a probability measure with a smaller support set, while preserving its moments up to a certain degree. This method is the subject of this paper and referred to as high-order recombination method. The aim of this paper is to present a theoretically guaranteed and practically feasible algorithm for the high-order recombination method, and to show that the method is applicable to real mathematical finance problems while maintaining the theoretically predicted performance.
1.1. Weak approximation problem of SDEs
Let $(\Omega,\mathcal{F},P)$ be a probability space, $\begin{pmatrix}B^{1}(t)&\dots&B^{d}(t)\end{pmatrix}$ be the $d$-dimensional standard Brownian motion, $B^{0}(t)=t$, and $C^{\infty}_{b}(\mathbb{R}^{N};\mathbb{R}^{N})$ be the set of $\mathbb{R}^{N}$-valued smooth functions defined in $\mathbb{R}^{N}$ whose derivatives of any order are bounded. $I_{N}$ denotes the identity map on $\mathbb{R}^{N}$. We also let $X(t,x)$ denote an $\mathbb{R}^{N}$-valued diffusion process defined by the Stratonovich-form SDE:
$X(t,x)=x+\sum_{i=0}^{d}\int_{0}^{t}V_{i}I_{N}(X(s,x))\circ\mathrm{d}B^{i}(s),$ (1)
where $x\in\mathbb{R}^{N}$ and $V_{0},\dots,V_{d}$ are tangent vector fields on $\mathbb{R}^{N}$ whose coefficients belong to $C^{\infty}_{b}(\mathbb{R}^{N};\mathbb{R}^{N})$, that is $V_{i}I_{N}=\begin{pmatrix}(V_{i}I_{N})^{1}&\dots&(V_{i}I_{N})^{N}\end{pmatrix}\in C_{b}^{\infty}(\mathbb{R}^{N};\mathbb{R}^{N})$, and the equality $V_{i}g(y)=\displaystyle{\sum_{j=1}^{N}(V_{i}I_{N})^{j}(x)\frac{\partial g}{\partial x_{j}}(y)}$ holds for all $i\in\{0,1,\dots,d\}$, $g\in C^{\infty}(\mathbb{R}^{N})$ and $y\in\mathbb{R}^{N}$. Furthermore, SDE (1) can be expressed in an alternative form by utilising the Ito integral $\mathrm{d}B^{i}(t)$:
$X(t,x)=x+\int_{0}^{t}\tilde{V_{0}}(X(s,x))\,\mathrm{d}t+\sum_{i=1}^{d}\int_{0}^{t}V_{i}I_{N}(X(s,x))\,\mathrm{d}B^{i}(s),$ (2)
where
$(\tilde{V}_{0}I_{N})^{k}=\left(V_{0}I_{N}\right)^{k}+\frac{1}{2}\sum_{i=1}^{d}V_{i}\left(V_{i}I_{N}\right)^{k}$ (3)
for all $k\in\{1,\dots,N\}$. This equality is called Ito–Stratonovich transformation.
In this paper, we address the weak approximation problem of SDEs, that is the numerical calculation of $E[f\left(X\left(T,x\right)\right)]$, where $X\left(t,x\right)$ is the value of an $N$-dimensional diffusion process at time $t$ and $f$ is a $\mathbb{R}$-valued function defined on $\mathbb{R}^{N}$ with some regularity. This calculation is crucial in many areas of applied science and engineering, for example, pricing and hedging financial derivatives. There are two main approaches to solve this problem numerically: using numerical methods for partial differential equations and performing simulations. Here, we focus on simulations, particularly on so-called tree-based simulation method described in the next subsection.
1.2. High-order cubature on Wiener space method
We adopt the high-order cubature on Wiener space methods [18, 22] in this paper. In the cubature on Wiener space methods, one approximates the infinite-dimensional Wiener measure with a finite-dimensional measure on a given partition of time $[0,T]$. We refer to this partition as $\Delta=\{0=t_{0}<\dots<t_{n}=T\}$ and, define $\sharp\Delta=n$ and $s_{j}=t_{j}-t_{j-1}$.
1.2.1. cub3: a cubature of degree $3$
As a simple illustrative example of cubature measure on Wiener space, a cubature of degree 3, called cub3, which is proposed in [18, 24] is outlined below.
$\begin{split}X^{({\rm cub3},\Delta)}_{t_{0}}(\eta)&=x,\\ X^{(\rm{cub3},\Delta)}_{\lambda t_{j+1}+(1-\lambda)t_{j}}(\eta)&=\exp\left(\lambda\left(s_{j+1}V_{0}+\sum_{i=1}^{d}\sqrt{s_{j+1}}\eta_{i,j+1}V_{i}\right)\right)X^{(\rm{cub3},\Delta)}_{t_{j}}(\eta)\end{split}$ (4)
where $0\leq\lambda\leq 1$, $\exp(V)x$ denotes the solution at time $1$ of the ordinary differential equation (ODE)
$\frac{\mathrm{d}z_{t}}{\mathrm{d}t}=VI_{N}\left(z_{t}\right),\;\;z_{0}=x,$ (5)
for a tangent vector field $V$ on $\mathbb{R}^{N}$ and $\eta=\left\{\eta_{i,j}\right\}_{\begin{subarray}{c}i=1,\dots,d\\ j=1,\dots,\sharp\Delta\end{subarray}}$ denotes a $d\sharp\Delta$-dimensional vector consisting of i.i.d. random variables such that $P(\eta_{i,j}=\pm 1)=1/2$. We remark that for $s\in\mathbb{R}$, $\exp(sV)(x)$ is well-defined by virtue of the chain rule and solves the ODE (5).
Then we obtain a $3$rd-order cubature measure on Wiener space $\mu^{(\rm{cub3},\Delta)}$ supported by $\left\{\omega_{\eta}\left|\,\eta\in\{-1,1\}^{d\sharp\Delta}\right.\right\}$ where
$\omega_{\eta}=\left(X_{s}^{(\rm{cub3},\Delta)}(\eta)\right)_{s\in[0,T]}$
and each $\omega_{\eta}$ has an equal weight of $2^{-d\sharp\Delta}$. Therefore, the computational cost of calculating
$\begin{split}\mu_{T}^{(\rm{cub3},\Delta)}(f)&=E\left[f\left(X^{(\rm{cub3},\Delta)}_{T}\right)\right]\\ &=\sum_{\eta\in\{-1,1\}^{d\sharp\Delta}}2^{-d\sharp\Delta}f\left(\left.\omega_{\eta}\right|_{s=T}\right)\end{split}$
grows exponentially with $\sharp\Delta$.
1.2.2. Tree-based simulation
Discretisation methods similar to the cub3 method, which proceeds by drawing a direct product of discrete random variables $\eta$, are called tree based simulations. This terminology stems from the structural interpretation that the support $\left\{\left.\omega_{\eta}\,\right|\,\eta\in\{-1,1\}^{d\sharp\Delta}\right\}$ of the measure can be considered as a tree, with the root $\left.\omega_{\eta}\right|_{s=0}$ and the leaves $\left\{\left.\left.\omega_{\eta}\right|_{s=T}\,\right|\,\eta\in\{-1,1\}^{d\sharp\Delta}\right\}$.
We will prepare some notations for tree-based simulations. Let $\{X^{({\rm Alg},\Delta)}\left(t_{i}\right)\}_{i=0,1,\dots,\sharp\Delta}$ denote a discretised process by a discretisation method Alg using discrete random variables with respect to the partition $\Delta$ of interval $[0,T]$. We denote a probability measure on $\mathbb{R}^{N}$ induced by $X^{({\rm Alg},\Delta)}\left(t_{i}\right)$ as $\mu_{t_{i}}^{({\rm Alg},\Delta)}=\sum_{j=1}^{m_{i}}w_{j}^{(i)}\delta_{p_{j}^{(i)}}$, where $\left\{\left(p_{j}^{(i)},w_{j}^{(i)}\right)\right\}_{j=1\dots,m_{i}}$ is the set of pairs consisting of points $p_{j}^{(i)}$ in $\mathbb{R}^{N}$ and weights $w_{j}^{(i)}$ assigned to them. $\delta_{x}$ denote the Dirac measure at $x\in\mathbb{R}^{N}$. We also denote the transition probability as
$w^{(i)}(k,j)=P\left(X^{({\rm Alg},\Delta)}\left(t_{i+1}\right)=p_{j}^{(i+1)}\left|\,X^{({\rm Alg},\Delta)}\left(t_{i}\right)=p_{k}^{(i)}\right.\right).$ (6)
Then, we define one step forward operator on the measure space as
$\hat{Q}^{\left({\rm Alg},\Delta\right)}_{(s_{i+1})}\left(\sum_{j=1}^{m_{i}}\alpha_{j}\delta_{p_{j}^{(i)}}\right)=\sum_{k=1}^{m_{i}}\sum_{j=1}^{m_{i+1}}\alpha_{k}w^{(i)}(k,j)\delta_{p_{j}^{(i+1)}},$ (7)
where $\alpha_{i}\in\mathbb{R}$. We also define
$w_{j}^{(i+1)}=\begin{cases}P\left(X^{(\rm{Alg},\Delta)}(t_{1})=p^{(1)}_{j}\right)&\text{if $i=0$}\\ \displaystyle{\sum_{k=1}^{m_{i}}w^{(i)}(k,j)w_{j}^{(i)}}&\text{if $1\leq i\leq\sharp\Delta-1$.}\end{cases}$ (8)
Then, we get the measure $\mu_{T}^{(\rm{Alg},\Delta)}=\hat{Q}^{\left({\rm Alg},\Delta\right)}_{(s_{\sharp\Delta})}\circ\cdots\circ\hat{Q}^{\left({\rm Alg},\Delta\right)}_{(s_{1})}\left(\delta_{x}\right)$. The process of approximating $E[f\left(X\left(T,x\right)\right)]$ by the expectation
$\mu_{T}^{(\rm{Alg},\Delta)}(f)=\sum_{j=1}^{m_{\sharp\Delta}}w_{j}^{(\sharp\Delta)}f\left(p_{j}^{(\sharp\Delta)}\right),$ (9)
is referred to as weak approximation by tree-based simulation. The cub 3 method $\mu_{T}^{(\rm{cub3},\Delta)}$, introduced in Subsection 1.2.1, is one of the simplest examples of tree-based simulation methods. The Ninomiya–Victoir method ([24]) $\mu_{T}^{(\rm{NV},\Delta)}$, which serves as the subject of the numerical experiments in this paper, is introduced in Subsection 2.3.
1.2.3. Order of discretisation
The order of discretisation method is defined as follows.
Definition 1.
We call Alg a $p$th order discretisation or discretisation of order $p$ if there exists a positive constant $C_{p}$ and for all $n\in\mathbb{N}$ a partitioning $\Delta$ of $[0,T]$ such that $\sharp\Delta=n$ exists and the following inequality
$\left\lvert E\left[f\left(X(t,x)\right)\right]-E\left[f\left(X^{({\rm Alg},\Delta)}\left(T\right)\right)\right]\right\rvert<C_{p}(\sharp\Delta)^{-p}$ (10)
holds.
In the following, we will refer to discretisation methods of second or higher order as high-order discretisation. Some specific methods of high-order discretisation method based on cubature on Wiener space are proposed in [24, 20, 26], which is introduced later in subsection 2.3. In addition, research on the construction of high-order cubature measures has explored various avenues, including simple algebraic construction [8], randomized construction [9], and Hopf algebras and unshuffle expansions-based construction [6]. For ODEs, a series of lower-order discretisation methods that are combined by allowing negative weights to obtain high-order discretisations are called extrapolations. With regard to the weak approximation of SDEs, the extrapolation of the Euler–Maruyama method is studied in [1]. Furthermore, as demonstrated in [7, 25], extrapolation methods can be integrated with [24] to derive discretisations of any order.
1.3. High-order recombination
When implementing the high-order cubature measure on Wiener space, one often suffers from the problem of exponential growth in the number of $\operatorname{supp}\left(\mu_{T}^{(\rm{Alg},\Delta)}\right)$ as the number of time partitions $\sharp\Delta$ increases. For example, it can be seen from subsection 1.2.1 that the cardinality of the support of $\mu_{T}^{(\rm{cub3},\Delta)}$ is $2^{d\sharp\Delta}$, while the cardinality of the support of $\mu_{T}^{(\rm{NV},\Delta)}$ is $9^{d\sharp\Delta}$, where NV is the second order discretisation method to be explained in subsection 2.3. In the following, we refer to this problem as the support explosion problem.
At present, the most widely practised methods for overcoming this problem are those of a partial sampling method, such as the Monte Carlo and quasi-Monte Carlo methods. Among the partial sampling methods, the method called TBBA is particularly applicable and effective for the tree-based simulation methods [2, 21, 19]. Furthermore, in [3, 4, 22], the authors successfully apply tree-based simulation with TBBA to the numerical calculations of forward backward SDEs.
High-order recombination is proposed in [17] as an alternative solution to the support explosion problem. There, the authors show that high-order recombination can be used to keep the growth rate of the cardinality of the support of the measure in polynomial order with respect to the required approximation accuracy. Several applications of high-order recombination are discussed in [15]. In [27], a number of algorithms for obtaining reduced measures are discussed. The notion of reduced measure plays a significant role in the construction of recombination measures, and will be introduced in the subsequent subsection 1.3.1.
An overview of the procedure for constructing what we call high-order recombination measures, which are the main subject of this paper, is given in this subsection.
1.3.1. Reduced measure
We first introduce the concept of a reduced measure, which is a measure with reduced support preserving relations defined by a set of test functions. Let $\mu$ be a discrete probability measure on a space $\tilde{\Omega}$, and $G=\{g_{1},g_{2},\dots,g_{\mathcal{M}_{G}}\}$ be a set of test functions, where $\mathcal{M}_{G}$ is the number of test functions and each $g_{i}:\tilde{\Omega}\rightarrow\mathbb{R}$ is a measurable function.
Definition 2.
(Reduced measure) A discrete probability measure $\tilde{\mu}$ is said to be a reduced measure from $\mu$ with respect to $G$ if it satisfies the following three conditions:
  1. (i)
    $\operatorname{supp}\left(\tilde{\mu}\right)\subset\operatorname{supp}\left(\mu\right)$,
  2. (ii)
    $\int g(x)\tilde{\mu}(dx)=\int g(x)\mu(dx)\quad\text{for all}\quad g\in G$,
  3. (iii)
    $\operatorname{card}\left(\operatorname{supp}\left(\tilde{\mu}\right)\right)\leq\mathcal{M}_{G}+1$,
where $\operatorname{card}(S)$ denotes the cardinality of the set $S$. $\operatorname{red}_{G}(\mu)$ denotes the set of reduced measures from $\mu$ with respect to $G$.
In [17], they show the existence of a reduced measure for a finite Borel measure $\mu$ on a polish space $\tilde{\Omega}$. In this paper, for the discretisation method of order $m$ and $\tilde{\Omega}=\mathbb{R}^{N}$, we consider the set of test functions $G$ as the distinct homogeneous polynomials of degree at most $m$ and assume that $g_{1}(x)=1$, $g_{2}(0)=g_{3}(0)=\dots=g_{\mathcal{M}_{G}}(0)=0$. Note that the condition (ii) determined by $g_{1}(x)=1$ ensures that the reduced measure $\tilde{\mu}$ also becomes a probability measure.
1.3.2. High-order recombination measures and their construction procedures
The high-order recombination measures $\left\{\mu^{(\rm{Alg},\Delta,\rm{Rec})}_{t_{i}}\right\}_{i=0}^{\sharp\Delta}$ in tree-based simulation with respect to $({\bf Alg},\Delta)$ are constructed recursively as follows:
$\begin{split}\mu^{(\rm{Alg},\Delta,\rm{Rec})}_{t_{0}}&=\delta_{x_{0}}\\ \nu^{(i)}&=\hat{Q}_{(s_{i})}^{(\rm{Alg},\Delta)}\left(\mu_{t_{i-1}}^{(\rm{Alg},\Delta,\rm{Rec})}\right)\\ \mu^{(\rm{Alg},\Delta,\rm{Rec})}_{t_{i}}&=\sum_{k=1}^{l_{i}}\nu^{(i)}\left(U^{(i)}_{k}\right)\tilde{\nu}^{(i)}_{k},\end{split}$ (11)
where $\left\{U^{(i)}_{k}\right\}_{k=1}^{l_{i}}$ is a set of subsets of $\operatorname{supp}\left(\nu^{(i)}\right)$ that satisfies
$\bigcup_{k=1}^{l_{i}}U^{(i)}_{k}=\operatorname{supp}\left(\nu^{(i)}\right)\quad\text{and}\quad U^{(i)}_{j}\cap U^{(i)}_{k}=\emptyset\quad\text{for all $j\neq k$,}$ (12)
$\tilde{\nu}^{(i)}_{k}\in\operatorname{red}_{G}\left(\left.\nu^{(i)}\right|_{U^{(i)}_{k}}\right)\quad\text{for $1\leq k\leq l_{i}$}$ (13)
and
$\left.\nu^{(i)}\right|_{U_{k}^{(i)}}=\frac{1}{\nu^{(i)}\left(U_{k}^{(i)}\right)}\sum_{p_{j}^{(i)}\in U_{k}^{(i)}}\hat{w}_{j}^{(i)}\delta_{p_{j}^{(i)}},$ (14)
which is the induced probability measure on $U^{(i)}_{k}$ from $\nu^{(i)}=\sum_{j=1}^{m_{i}}\hat{w}_{j}^{(i)}\delta_{p^{(i)}_{j}}$. Henceforth, a set $\left\{U_{k}\right\}_{k=1}^{l}$ that satisfies
$\bigcup_{k=1}^{l}U_{k}=\operatorname{supp}(\nu^{(i)})\quad\text{and}\quad U_{j}\cap U_{k}=\emptyset\quad\text{for all $j\neq k$}$ (15)
is referred to as a patch division of $\operatorname{supp}(\nu^{(i)})$. Note that this procedure is interwoven with the tree-based simulation procedure and is performed simultaneously. Also note that the total weights assigned to each patch are invariant throughout the recombination procedure. We therefore refer to this as the weight of the patch $U_{k}^{(i)}$ and denote it by $\operatorname{wt}\left(U^{(i)}_{k}\right)$, that is
$\operatorname{wt}\left(U^{(i)}_{k}\right)=\nu^{(i)}\left(U^{(i)}_{k}\right)=\mu_{t_{i}}^{(\rm{Alg},\Delta,\rm{Rec})}\left(U^{(i)}_{k}\right).$ (16)
A simplified illustration of these procedures can be found in Figure 1.
Figure 1. Illustration of recombination method
In [17] they show that we can proceed with the high-order recombination while maintaining the order of discretisation by choosing appropriate patch radii
$\operatorname{rad}\left(U_{j}\right)=\max\left\{\left\|p_{k}-p_{k^{\prime}}\right\|\left|\,p_{k},p_{k^{\prime}}\in U_{j}\right.\right\}.$ (17)
We can find numerical examples of the high-order recombination method for practical problems only in [23]. In the paper they show numerically that even with ad hoc patch division, which is not proved to be free from the support explosion problem, the recombination method can still be effective for certain problems.
In this paper, we provide a detailed explanation of the algorithms corresponding to the aforementioned procedure. Specifically, the explicit construction of $\hat{Q}_{(s_{i})}^{(\rm{NV},\Delta)}$ is presented in Subsection 2.3, an algorithm for the patch division of $\operatorname{supp}(\nu^{(i)})$ into $\left\{U^{(i)}_{k}\right\}_{k=1}^{l_{i}}$ is proposed in Subsection 3.2, and the algorithm for finding $\tilde{\nu}^{(i)}_{k}\in\operatorname{red}_{G}\left(\left.\nu^{(i)}\right|_{U^{(i)}_{k}}\right)$, which is the same algorithm used in [23], is provided in Appendix 1.
1.4. Contribution of this paper
This paper presents a novel recursive patch division algorithm for high-order recombination that addresses the support explosion problem, making it practically feasible for real-world applications in mathematical finance. Our contributions can be summarized as follows
  • An improved version of the error estimation given in [17] is presented, which considers the patch weights $\operatorname{wt}\left(U^{(i)}_{k}\right)$.
  • A recursive algorithm is proposed, which divides the support set into patches, taking into account both the radii and the weights of the patches. The algorithm contains a control parameter that can be varied to change the degree of recombination. The recursive nature of this algorithm allows for efficient and adaptive patch division based on the refined error evaluation.
  • An example of the application of this algorithm to some real financial option pricing problems is presented. In this example, the algorithm successfully avoids the support explosion problem as expected. It also shows how changing the parameter controlling the degree of recombination works in this example. n
The remainder of this paper is organized as follows. In Section 2, we introduce the necessary notation and preliminary results on high-order discretisation schemes and high-order recombination. In Section 3, we present the refined error evaluation and the novel recursive patch division algorithm, which are the main contributions of this paper. In Section 4, we provide numerical results to demonstrate the performance of our algorithm.
2. Preliminary
In this section, we introduce some notations and previous studies on high-order discretisation schemes and the recombination algorithm.
2.1. Free Lie algebras
Let $A=\{0,1,\dots,d\}$ and define the set of all words consisting of elements of $A$ as $A^{+}=\bigcup_{k=1}^{\infty}A^{k}$ and also define the set $A^{\ast}=\{\emptyset\}\cup A^{+}$. $A^{\ast}$ becomes a semigroup with identity $\emptyset$ when endowed with the concatenating product $\cdot$’, that is for arbitrary elements $\left(a_{1},\dots,a_{m}\right)$ and $\left(b_{1},\dots,b_{n}\right)$ of $A^{\ast}$,
$\left(a_{1},\dots,a_{m}\right)\cdot\left(b_{1},\dots,b_{n}\right)=\left(a_{1},\dots,a_{m},b_{1},\dots,b_{n}\right)$
and
$\emptyset\cdot\left(a_{1},\dots,a_{m}\right)=\left(a_{1},\dots,a_{m}\right)\cdot\emptyset=\left(a_{1},\dots,a_{m}\right).$
To simplify the notation, we identify $(a_{1},a_{2},\ldots,a_{m})\in A^{\ast}$ with $a_{1}\cdot a_{2}\cdots a_{m}$. If there is no risk of confusion, $\cdot$’ is also omitted. That is, we would rather write $a_{1}a_{2}\ldots a_{m}$ instead of $(a_{1},a_{2},\ldots,a_{m})$ or $a_{1}\cdot a_{2}\cdots\cdot a_{m}$. For $\alpha=a_{1}a_{2}\dots a_{k}\in A^{\ast}$, we define the length of $\alpha$ as $\lvert\alpha\rvert=k$ and $\lvert\emptyset\rvert=0$. We also define the order of $\alpha$ as $\|\alpha\|=|\alpha|+\operatorname{card}\left(\{1\leq i\leq k\;|\;a_{i}=0\}\right)$. Let $A^{\ast}_{1}=A^{\ast}\backslash\{\emptyset,0\}$ and $A^{\ast}(m)=\{\alpha\in A^{\ast}\;|\;\|\alpha\|\leq m\}$ and $A^{\ast}_{1}(m)=\{\alpha\in A^{\ast}_{1}\;|\;\|\alpha\|\leq m\}$ respectively for $m\in\mathbb{N}$.
Let $\mathbb{R}\langle A\rangle$ be the free $\mathbb{R}$-algebra with basis $A^{\ast}$ and $\mathbb{R}\langle\langle A\rangle\rangle$ be its completion, that is $\mathbb{R}\langle A\rangle=\left\{\left.\sum_{k=1}^{l}r_{k}\alpha_{k}\;\right|\;l\in\mathbb{N},r_{k}\in\mathbb{R},\alpha_{k}\in A^{\ast}\right\}$ and $\mathbb{R}\langle\langle A\rangle\rangle=\left\{\left.\sum_{\alpha\in A^{\ast}}r_{\alpha}\alpha\;\right|\;r_{\alpha}\in\mathbb{R}\right\}$. The latter is the set of all $\mathbb{R}$-coefficient formal series with basis $A^{\ast}$. For $R=\sum_{\alpha\in A^{\ast}}r_{\alpha}\alpha$ and $S=\sum_{\beta\in A^{\ast}}s_{\beta}\beta\in\mathbb{R}\langle A\rangle$, let their product $RS$ be defined by
$RS=\sum_{\theta=\alpha\beta\in A^{\ast}}r_{\alpha}s_{\beta}\theta,$
and their inner product $\langle R,S\rangle$ by
$\langle R,S\rangle=\sum_{\alpha=\beta}r_{\alpha}s_{\beta}.$
We also define $\|R\|_{2}=\left(\langle R,R\rangle\right)^{1/2}$. The Lie bracket is defined to be $[R,S]=RS-SR$ as usual. ${\mathcal{L}}_{\mathbb{R}}(A)$ denotes the set of Lie polynomials in $\mathbb{R}\langle A\rangle$, that is the smallest $\mathbb{R}$-submodule of $\mathbb{R}\langle A\rangle$ including $A$ and closed under the Lie bracket.
Let $\Phi$ be the homomorphism from $\mathbb{R}\langle A\rangle$ to the $\mathbb{R}$-algebra consisting of smooth differential operators over $\mathbb{R}^{N}$ such that
$\begin{split}\Phi(\emptyset)&=\text{Id},\\ \Phi(\alpha)&=V_{a_{1}}V_{a_{2}}\cdots V_{a_{k}}\mbox{\hskip 5.69054pt for \hskip 5.69054pt}\alpha=a_{1}a_{2}\dots a_{k}\in A^{+}.\end{split}$ (18)
Here Id is the identity operator.
We remark that $\Phi$ induces canonically a Lie algebra homomorphism from $\mathcal{L}_{\mathbb{R}}(A)$ to the Lie algebra of vector fields generated by $\left\{V_{0},\dots,V_{d}\right\}$. We denote $V_{R}=\Phi(R)$ for $R\in\mathbb{R}\langle A\rangle$ and $[\alpha]=[[\dots,[a_{1},a_{2}]\dots],a_{k}]$ for $\alpha=a_{1}a_{2}\dots a_{k}\in A^{\ast}$.
We remark that the following equality
$f(\exp(t\Phi(R))(x)=\Phi\left(j_{m}\left(\exp\left(tR\right)\right)\right)f(x)+O(t^{m+1})$ (19)
holds for $R\in{\mathcal{L}}_{\mathbb{R}}(A)$ by virtue of the Taylor expansion formula.
2.2. High-order discretisation scheme
In this subsection, we introduce high-order discretisation schemes of SDEs and related results. $C_{\rm{Lip}}(\mathbb{R}^{N};\mathbb{R})$ denotes the space of Lipschitz functions defined on $\mathbb{R}^{N}$, $\|\cdot\|_{\rm{Lip}}$ denotes the Lipschitz norm, that is
$\|f\|_{\rm{Lip}}=\sup_{x,y\in\mathbb{R}^{N},x\neq y}\frac{\left|f(x)-f(y)\right|}{\left|x-y\right|},$
and $\|\cdot\|_{\infty}$ denotes the uniform norm, that is $\left\|f\right\|_{\infty}=\sup_{x\in\mathbb{R}^{N}}\left|f(x)\right|$ for $f\in C_{\rm{Lip}}(\mathbb{R}^{N};\mathbb{R})$.
Definition 3 ($\left\{P^{X}_{t}\right\}_{t\geq 0}$).
Let $X(t,x)$ be the $\mathbb{R}^{N}$-valued diffusion process defined by (1). For $f\in C_{\rm{Lip}}(\mathbb{R}^{N};\mathbb{R})$ and $t\geq 0$, we define
$(P^{X}_{t}f)(x)=E\left[f(X(t,x))\right]$
and obtain a semigroup of linear operators $\{P^{X}_{t}\}_{t\geq 0}$ on $C_{\rm{Lip}}(\mathbb{R}^{N};\mathbb{R})$.
At first, we present an approximation theorem for high-order discretisation scheme, based on [11, 12] and [13].
Definition 4 (UFG condition [12]).
Let $\ell$ be an integer. The vector fields $V_{0},V_{1},\dots,V_{d}$ are said to satisfy the UFG condition, if for all $\alpha\in A^{\ast}_{1}$, there exists $\left\{{\varphi}_{{\alpha},{\beta}}\>\left|\;\beta\in A^{\ast}_{1}\left(\ell\right)\right.\right\}\subset C^{\infty}_{b}(\mathbb{R}^{N};\mathbb{R})$ satisfying
$V_{[\alpha]}=\sum_{\beta}\varphi_{{\alpha},{\beta}}V_{[\beta]}.$
Hereinafter, $\ell$ is referred to as the degree of the UFG condition. Then, we define the condition on approximation operators as follows.
Theorem 1 (Smoothness of the semigroup $\left\{P^{X}_{t}\right\}_{t\geq 0}$ in certain directions [12]).
Assume that the vector fields $V_{0},V_{1},\dots,V_{d}$ satisfy the UFG condition and $X$ is the diffusion process defined as (1). Let $\alpha_{1},\alpha_{2},\dots,\alpha_{k}\in A^{\ast}_{1}$. Then, there exists a constant $C>0$ such that
$\left\|V_{[\alpha_{1}]}\cdots V_{[\alpha_{k^{\prime}}]}P^{X}_{t}V_{[\alpha_{k^{\prime}+1}]}\cdots V_{[\alpha_{k}]}f\right\|_{L^{p}(dx)}\leq Ct^{\left(\|\alpha_{1}\|+\cdots+\|\alpha_{k}\|\right)/2}\|f\|_{L^{p}(dx)}$
Definition 5 ($m$-similar operator).
Markov operator $\tilde{Q}_{(s)}$ is said to be an $m$-similar operator if there exist ${\mathcal{L}}_{\mathbb{R}}(A)$-valued random variables $L_{(s)}^{(1)},L_{(s)}^{(2)},\dots,L_{(s)}^{(K)}$ such that
$E\left[j_{m}\left(\exp\left(L_{(s)}^{(1)}\right)\exp\left(L_{(s)}^{(2)}\right)\cdots\exp\left(L_{(s)}^{(K)}\right)\right)\right]=E\left[j_{m}\left(\sum_{\alpha\in A^{\ast}}B^{\circ\alpha}(s)\alpha\right)\right],$
$E\left[\left|\left|j_{m}\left(L_{(s)}^{(k)}\right)\right|\right|_{2}^{n}\right]<\infty,\quad\left\langle L_{(s)}^{(k)},(0)\right\rangle=1\quad\text{for $n\in\mathbb{N}$, $k=1,2,\dots K$},$
and
$\left(\tilde{Q}_{(s)}f\right)(x)\\ =E\left[f\left(\left(\exp\left(\Phi\left(L_{(s)}^{(1)}\right)\right)\exp\left(\Phi\left(L_{(s)}^{(2)}\right)\right)\cdots\exp\left(\Phi\left(L_{(s)}^{(K)}\right)\right)\right)(x)\right)\right].$
Here $B^{\circ\alpha}(s)$ is defined for $\alpha\in A^{*}$ as follows:
$B^{\circ\alpha}(t)=\begin{cases}1&\text{if $\alpha=\emptyset$},\\ \displaystyle{\int_{0}^{t}B^{\alpha^{\prime}}(s)\,\circ\mathrm{d}B^{a}(s)}&\text{if $\alpha=\alpha^{\prime}a$ where $\alpha^{\prime}\in A^{*}$ and $a\in A$}.\end{cases}$
The $m$-similar operator is intuitively understood as an approximation of $E[f\left(X\left(T,x\right)\right)]$ up to the $m$th order using a stochastic Taylor expansion of the solution of the SDE (1). We refer to $m$ as the degree of the discretisation method. Note that this definition has been slightly extended from the original definition used in [11, 12, 13].
Theorem 2 (Approximation theorem for high-order discretisation schemes [13]).
Assume that the vector fields of the SDE (1) satisfy the UFG condition, $X$ is the diffusion process defined by (1) and $Q_{(s)}$ is an $m$-similar operator. Then there exist positive constants $C$ and $C^{\prime}$ which depend only on $m$ and the vector fields $V_{0},V_{1},\dots,V_{d}$ and satisfy
$\left\|P^{X}_{s}f-Q_{(s)}f\right\|_{\infty}\leq Cs^{(m+1)/{2}}\left\|f\right\|_{\rm{Lip}}$
and
$\left\|Q_{(s)}f\right\|_{\infty}\leq\exp(C^{\prime}s)\left\|f\right\|_{\infty}$
for all $f\in C_{\rm{Lip}}(\mathbb{R}^{N};\mathbb{R})$.
Consequently, the total discretisation is bounded as follows. Let $T>0$ be a fixed time and consider a partition $\Delta$ of interval $[0,T]$ such that $\sharp\Delta=n$ as
$t_{0}=0<t_{1}<\dots<t_{n}=T.$
Then there exists a constant $C^{\prime\prime}$ depending only on $m$, the vector fields $V_{0},V_{1},\dots,V_{d}$, and $T$ such that
$\left\|P^{X}_{T}f-Q_{(s_{n})}\circ Q_{(s_{n-1})}\circ\cdots\circ Q_{(s_{1})}f\right\|_{\infty}\\ \leq C^{\prime\prime}\left\|f\right\|_{\rm{Lip}}\left(s_{n}^{1/2}+\sum_{j=m}^{m+1}\sum_{i=1}^{n-1}\frac{s_{i}^{(j+1)/2}}{(T-t_{i})^{j/2}}\right),$ (20)
where $s_{k}=t_{k}-t_{k-1}$.
Remark 1 (The regularity assumption on $f$).
In the initial work of Kusuoka [11, 13], Theorem 2 is proved on the assumption of $f\in C_{b}^{\infty}(\mathbb{R}^{N};\mathbb{R})$. It is easily seen that this assumption can be relaxed to $f\in C_{\rm{Lip}}(\mathbb{R}^{N};\mathbb{R})$, as pointed out in Section 3 of [18].
2.3. N–V method
We introduce the N–V method from [24], which is a second-order discretisation. We denote by $(G_{1}\circ G_{2})g(x)$ the composition $(G_{1}(G_{2}g))(x)$ of $G_{1}$ and $G_{2}:C_{\rm{Lip}}(\mathbb{R}^{N};\mathbb{R})\rightarrow C_{\rm{Lip}}(\mathbb{R}^{N};\mathbb{R})$ hereinafter.
Theorem 3 ([24, 14]).
Let $\eta_{1},\eta_{2},\dots,\eta_{d}$ be independent standard normal random variables and
$\begin{split}Q_{(s)}^{\rm{(NV})}f(x)=&E\biggl{[}\frac{1}{2}f\left(\exp\left(sV_{0}\right)\exp\left(\sqrt{s}\eta_{1}V_{1}\right)\cdots\exp\left(\sqrt{s}\eta_{d}V_{d}\right)(x)\right)\\ &+\frac{1}{2}f\left(\exp\left(\sqrt{s}\eta_{d}V_{d}\right)\cdots\exp\left(\sqrt{s}\eta_{1}V_{1}\right)\exp\left(sV_{0}\right)(x)\right)\biggr{]}.\end{split}$ (21)
Then $Q^{(\rm{NV})}_{(s)}$ are 5-similar operators and the N–V method $Q^{(\rm{NV})}_{(T/n)}\circ Q^{(\rm{NV})}_{(T/n)}\circ\cdots\circ Q^{(\rm{NV})}_{(T/n)}$ is a second order discretisation method.
Replacing the standard normal random variables ${\eta}$ with the discrete random variables $\hat{\eta}^{(\text{$5$th})}$, we define the discrete N–V method as in  [22]. Let $\hat{\eta}^{(\text{$5$th})}$ be a discrete random variables on $(\hat{\Omega},\hat{\mathcal{F}},\hat{P})$ defined as
$\hat{P}\left(\hat{\eta}^{(\text{$5$th})}=\pm\sqrt{3}\right)=\frac{1}{6},\;\;\hat{P}\left(\hat{\eta}^{(\text{$5$th})}=0\right)=\frac{2}{3}.$ (22)
We note that the moments of $\hat{\eta}^{(\text{$5$th})}$ up to the $5$th order are equal to those of the standard normal random variables. As is subsection 1.2.2, we denote a probability measure on $\mathbb{R}^{N}$ induced by $X^{({\rm Alg},\Delta)}\left(t_{i}\right)$ as $\mu_{t_{i}}^{({\rm Alg},\Delta)}=\sum_{j=1}^{m_{i}}w_{j}^{(i)}\delta_{p_{j}^{(i)}}$, where $\left\{\left(p_{j}^{(i)},w_{j}^{(i)}\right)\right\}_{j=1\dots,m_{i}}$ is the set of pairs consisting of points $p_{j}^{(i)}$ in $\mathbb{R}^{N}$ and weights $w_{j}^{(i)}$ assigned to them. Let $\{\hat{\eta}^{(\text{$5$th})}_{i,l}\}_{\begin{subarray}{c}i=1,\dots,n\\ l=1,\dots,d\end{subarray}}$ be a set of independent and identically distributed discrete random variables following the same distribution as $\hat{\eta}^{(\text{$5$th})}$. Let $\phi$ be a function from {±$\{\pm\sqrt{3},0\}^{d}$ to $[0,1]$ defined by
$\phi(\eta)=\left(\frac{1}{6}\right)^{\operatorname{card}\{i\;|\;\eta_{i}=\pm\sqrt{3}\}}\left(\frac{2}{3}\right)^{\operatorname{card}\{i\;|\;\eta_{i}=0\}}$
for $\eta=\left(\eta_{1},\dots,\eta_{d}\right)\in\{\pm\sqrt{3},0\}^{d}$. We also define functions $y^{(+)},y^{(-)}$ from $\mathbb{R}^{N}\times\{\pm\sqrt{3},0\}^{d}\times\mathbb{R}_{>0}$ to $\mathbb{R}^{N}$ as
$\begin{split}y^{(+)}\left(x,\eta,s\right)&=\exp\left(sV_{0}\right)\exp\left(\sqrt{s}\eta_{1}V_{1}\right)\cdots\exp\left(\sqrt{s}\eta_{d}V_{d}\right)\left(x\right),\\ y^{(-)}\left(x,\eta,s\right)&=\exp\left(\sqrt{s}\eta_{d}V_{d}\right)\cdots\exp\left(\sqrt{s}\eta_{1}V_{1}\right)\exp\left(sV_{0}\right)\left(x\right).\end{split}$
Then, the operator $\hat{Q}_{(s_{i})}^{({\rm NV},\Delta)}$ between discrete measures can be defined as
$\hat{Q}_{(s_{i})}^{({\rm NV},\Delta)}\left(\sum_{j=1}^{m_{i}}w_{j}^{(i)}\delta_{p_{j}^{(i)}}\right)\\ =\sum_{j=1}^{m_{i}}\sum_{\eta\in\{\pm\sqrt{3},0\}^{d}}w_{j}^{(i)}\frac{\phi(\eta)}{2}\left(\delta_{y^{(+)}(p_{j}^{(i)},\eta,s_{i})}+\delta_{y^{(-)}(p_{j}^{(i)},\eta,s_{i})}\right).$ (23)
In this way, a concrete example of $\hat{Q}^{\left({\rm Alg},\Delta\right)}_{(s_{i})}$ used in this paper, as discussed in subsection 1.2.2, is constructed. We apply the recombination algorithm described in subsection 1.3.2 to this operator and obtain $\mu_{T}^{\left({\rm NV},\Delta,{\rm Rec}\right)}$.
Similarly, for an $m$-moment similar operator $Q_{(s)}$, we define $\hat{Q}_{(s_{i})}(\mu)$ by adapting discrete the random variables $\hat{\eta}^{(m\text{th})}$ whose moments up to the $m$th order are matched with those of the standard normal random variables. In this paper, we consider a $p$th order discretisation algorithm defined from $p$-moment similar operator and its recombination measure.
Remark 2 (The number of random variables involved in the N–V method).
In [24], the authors propose to implement the N–V method using the Bernoulli random variables in addition to the standard normal random variables $\eta_{1},\eta_{2},\dots,\eta_{d}$, in order to halve the number of ODEs to be solved. In contrast, the present paper employs the N-V method without utilising Bernoulli random variables, instead calculating the expectation as shown in equation (21). This approach is motivated by the necessity to minimise the number of support points of the discrete N–V method in order to regulate the overall calculation cost, as noted in Remark 3.
Remark 3 (Comparison between high-order discretisation methods).
The N–V method has the following advantages compared to other high-order discretisation methods such as [20, 26]. As stated in the original paper [24], the N–V method is often more efficiently implementable, because analytical solutions for ODEs are available in some cases such as the Heston model [10]. Furthermore, the Feller condition [5] ensures the positivity of a process discretised by the N–V method.
Moreover, when we implement the high-order discretisation methods with discrete random variables, the number of random variables involved in the N–V method is parsimonious in the sense that only $d$ random numbers are required to approximate an SDE driven by the $d$-dimensional Brownian motion. The numbers of random variables and supports of the measures are summarized in Table 1.
Method
Num. $\{\hat{\eta}\}$
in $\hat{Q}_{(1/n)}$
Num. support of measure
of $\hat{Q}_{(T/n)}\circ\hat{Q}_{(T/n)}\circ\cdots\circ\hat{Q}_{(T/n)}$
N–V $d$ $(2\times 3^{d})^{n}$
N–N [20] $2d$ $3^{2dn}$
$Q_{(s)}^{(7,2)}$($d=2$) [26] 7 $5^{7n}$
Table 1. Number of supports of $\hat{X}^{(\text{method})}_{s,x}$, $\sharp\Delta=n$
2.4. High-order recombination: approximation theorem
In this subsection, we introduce the approximation theorem established by [17]
Definition 6 (UH condition).
The vector fields $V_{0},V_{1},\dots,V_{d}$ are said to satisfy the uniform Hörmandar (UH) condition, if there exists $\ell\in\mathbb{N}$ such that
$\inf_{x,y\in\mathbb{R}^{N},|y|=1}\left(\sum_{\alpha\in A^{\ast}_{1}(\ell)}\left\langle V_{[\alpha]}(x),y\right\rangle\right)>0$
Theorem 4 ([17] Theorem 19).
Assume that the vector fields $V_{0},V_{1},\dots,V_{d}$ satisfy the UH condition. Let Alg be an $m$th order discretisation and $\Delta$ be a partition of $[0,T]$ as $0=t_{0}<t_{1}<\dots<t_{\sharp\Delta}=T$. Let $\left\{\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)},U_{k}^{(i)}\right\}_{\begin{subarray}{c}i=0,\dots,\sharp\Delta,\\ k=1,\dots,l_{i}\end{subarray}}$ be a set of recombination measures and patch divisions of their supports. Then,
$\left\|P^{X}_{T}f-\mu_{T}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\left(f\right)\right\|_{\infty}\\ \leq\left(C_{1}(T)\left(s_{\sharp\Delta}^{1/2}+\sum_{i=1}^{\sharp\Delta-1}\sum_{j=m}^{m+1}\frac{s_{i}^{(j+1)/2}}{(T-t_{i})^{j/2}}\right)+C_{2}(T)\sum_{i=2}^{\sharp\Delta-1}\frac{u_{i}^{m+1}}{(T-t_{i})^{m\ell}}\right)\|f\|_{\rm{Lip}}$ (24)
holds where $u_{i}=\max\left.\left\{\operatorname{rad}\left(U_{k}^{(i)}\right)\;\right|\;k=1,\dots,l_{i}\right\}$ and $s_{i}=t_{i}-t_{i-1}$.
Example 1 (Patch condition ([17] Examples 21)).
Given a time step as
$t_{i}=T\left(1-\left(1-\frac{j}{n}\right)^{\gamma}\right)$
for some $\gamma>0$. If we set
$u_{i}=\left(\frac{s_{i}^{m+1}}{(T-t_{i})^{m-r\ell}}\right)^{1/(2(l+1))},$
then we have
$\left\|P^{X}_{T}f-\mu_{T}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\left(f\right)\right\|_{\infty}\leq\frac{C_{2}(T)}{(\sharp\Delta)^{(m-1)/2}}\|f\|_{\rm{Lip}}.$ (25)
In other word, the recombination do not harm the order of discretisation error.
3. Refined error estimation and patch division algorithms
In this section, we present the main results of this paper. First, we refine Theorem 4 and Example 1 by extending the argument in [17] to obtain Theorem 5 and an example of a sufficient condition (Example 2) for the patch division to satisfy the required recombination error.
3.1. Refined error estimation
Theorem 5.
Assume that the vector fields $V_{0},V_{1},\dots,V_{d}$ satisfy the UH condition. Let Alg be an $m$th order discretisation and $\Delta$ be a partition of $[0,T]$ as $0=t_{0}<t_{1}<\dots<t_{\sharp\Delta}=T$. Let $\left\{\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)},U_{k}^{(i)}\right\}_{\begin{subarray}{c}i=0,\dots,\sharp\Delta,\\ k=1,\dots,l_{i}\end{subarray}}$ be a set of recombination measures and patches divisions of their supports. Then,
$\begin{split}&\|P^{X}_{T}f-\mu_{T}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\left(f\right)\|_{\infty}\\ &\leq\Biggl{(}C_{1}(T)\left(s_{\sharp\Delta}^{1/2}+\sum_{i=1}^{\sharp\Delta-1}\sum_{q=m+1}^{K}\frac{s_{i}^{q/2}}{(T-t_{i})^{(q-1)/2}}\right)\\ &\qquad\qquad\qquad+C_{2}(T)\sum_{i=1}^{\sharp\Delta-1}\frac{\sum_{k=1}^{l_{i}}\left(\operatorname{rad}\left(U_{k}^{(i)}\right)\right)^{m+1}\operatorname{wt}\left(U_{k}^{(i)}\right)}{(T-t_{i})^{m/2}}\Biggr{)}\|\nabla f\|_{\infty}.\end{split}$ (26)
for some $Q>0$ and $s_{i}=t_{i}-t_{i-1}$.
Hereinafter, we refer to the first term of the right-hand side of (26)
$C_{1}(T)\left(s_{\sharp\Delta}^{1/2}+\sum_{i=1}^{\sharp\Delta-1}\sum_{q=m+1}^{K}\frac{s_{i}^{q/2}}{(T-t_{i})^{(q-1)/2}}\right)\|\nabla f\|_{\infty}$
as the discretisation error and the second term
$C_{2}(T)\sum_{i=1}^{\sharp\Delta-1}\frac{\sum_{k=1}^{l_{i}}\left(\operatorname{rad}\left(U_{k}^{(i)}\right)\right)^{m+1}\operatorname{wt}\left(U_{k}^{(i)}\right)}{(T-t_{i})^{m/2}}\|\nabla f\|_{\infty}$
as the recombination error.
Proof.
Let $\hat{Q}_{(s)}$ be an $m$-similar operator. From Theorem 3 of [13], there exists $Q\in\mathbb{N}$ such that
$\begin{split}\left\|P^{X}_{s}f-\hat{Q}_{(s)}f\right\|_{\infty}\leq c_{1}\left(s^{{(m+1)}/{2}}\|\nabla f\|_{\infty}+\sum_{q=m+1}^{K}s^{{q}/{2}}\|f\|_{V,q}\right)\end{split}$ (27)
and
$\begin{split}\left\|P^{X}_{s}f-\hat{Q}_{(s)}f\right\|_{\infty}\leq c_{3}s^{1/2}\|\nabla f\|_{\infty}\end{split}$ (28)
hold for $s>0$, where $\|\cdot\|_{V,q}$ is a semi-norm on $C_{b}^{\infty}\left(\mathbb{R}^{N};\mathbb{R}\right)$ defined as
$\|f\|_{V,q}=\sum_{q^{\prime}=1}^{q}\sum_{\alpha_{1}\cdots\alpha_{q^{\prime}}\in A^{\ast}}\left\|V_{[\alpha_{1}]}\cdots V_{[\alpha_{q^{\prime}}]}f\right\|_{\infty}.$ (29)
We also have the following estimation from Theorem 2 of [12].
$\begin{split}\left\|P^{X}_{T-t}f\right\|_{V,q}\leq c_{2}\frac{\|\nabla f\|_{\infty}}{\left(T-t\right)^{(q-1)/2}}.\end{split}$ (30)
Then, setting $\nu^{(i)}=\hat{Q}_{(s_{i})}^{(\rm{Alg},\Delta)}\left(\mu_{t_{i-1}}^{(\rm{Alg},\Delta,\rm{Rec})}\right)$ as in (11), we decompose the total error as follows. Note that the operator $P^{X}_{T}$ and the measure $\mu_{T}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}$, and its intermediate measures $\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}$, $\nu^{(i)}$ depend on the initial value $x_{0}$.
$\begin{split}&\sup_{x_{0}\in\mathbb{R}^{N}}\left|P^{X}_{T}f(x)-\mu_{T}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\left(f\right)\right|\\ &\qquad\qquad\quad\leq\left\|P^{X}_{T}f-\hat{Q}_{(s_{1})}P_{T-t_{1}}f\right\|_{\infty}\\ &\qquad\qquad\qquad\quad+\sum_{i=1}^{n-1}\left\|\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\left(P^{X}_{T-t_{i}}f\right)-\nu^{(i+1)}\left(P^{X}_{T-t_{i+1}}f\right)\right\|_{\infty}\\ &\qquad\qquad\qquad\quad+\sum_{i=1}^{n-1}\left\|\nu^{(i)}\left(P^{X}_{T-t_{i}}f\right)-\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\left(P^{X}_{T-t_{i}}f\right)\right\|_{\infty}.\end{split}$ (31)
The first and second terms of right-hand side of (31), which correspond to the discretisation error, can be estimated as follows.
$\begin{split}&\left\|P^{X}_{T}f-\mu_{x_{0},\Delta,\mathcal{U}}^{(1)}P^{X}_{T-t_{1}}f\right\|_{\infty}+\sum_{i=1}^{n-1}\left\|\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\left(P^{X}_{T-t_{i}}f\right)-\nu^{(i+1)}\left(P^{X}_{T-t_{i+1}}f\right)\right\|_{\infty}\\ &\;\;=\left\|P^{X}_{s_{1}}P^{X}_{T-t_{1}}f-\hat{Q}_{(s_{1})}P^{X}_{T-t_{1}}f\right\|_{\infty}+\sum_{i=1}^{n-1}\left\|E_{\tilde{\mu}_{x_{0},\Delta,\mathcal{U}}^{(i)}}\left(P^{X}_{s_{i+1}}-\hat{Q}_{(s_{i+1})}\right)P^{X}_{T-t_{i+1}}f\right\|_{\infty}\\ &\;\;\leq\left\|P^{X}_{s_{1}}P^{X}_{T-t_{1}}f-\hat{Q}_{(s_{1})}P^{X}_{T-t_{1}}f\right\|_{\infty}+\sum_{i=1}^{n-1}\left\|\left(P^{X}_{s_{i+1}}-\hat{Q}_{(s_{i+1})}\right)P^{X}_{T-t_{i+1}}f\right\|_{\infty}\end{split}$
Then, using (27), (28), (30), we have
$\begin{split}&\left\|P^{X}_{T}f-\hat{Q}_{(s_{1})}\left(P_{T-t_{1}}\right)\right\|_{\infty}+\sum_{i=1}^{n-1}\left\|\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\left(P^{X}_{T-t_{i}}f\right)-\nu^{(i+1)}\left(P^{X}_{T-t_{i+1}}f\right)\right\|_{\infty}\\ &\;\;\leq\sum_{i=1}^{n-1}c_{1}\left({s_{i}}^{{(m+1)}/{2}}\left\|\nabla P^{X}_{T-t_{i}}f\right\|_{\infty}+\sum_{q=m+1}^{K}{s_{i}}^{{q}/{2}}\left\|P^{X}_{T-t_{i}}f\right\|_{V,q}\right)+c_{3}{s_{n}}^{{1}/{2}}\|\nabla f\|_{\infty}\\ &\;\;\leq\sum_{i=1}^{n-1}c_{1}\left({s_{i}}^{{(m+1)}/{2}}\left\|\nabla P^{X}_{T-t_{i}}f\right\|_{\infty}+\sum_{q=m+1}^{K}{s_{i}}^{{q}/{2}}\frac{\left\|\nabla f\right\|_{\infty}}{\left(T-t_{i}\right)^{{q}/{2}}}\right)+c_{3}{s_{n}}^{{1}/{2}}\|\nabla f\|_{\infty}\\ &\;\;\leq C_{1}(T)\left({s_{n}}^{1/2}+\sum_{i=1}^{n-1}\sum_{q=m+1}^{K}\frac{{s_{i}}^{q/2}}{(T-t_{i})^{(q-1)/2}}\right)\|\nabla f\|_{\infty}\end{split}$
The third term of (31) corresponds to the recombination error and can be bound by using Corollary 14 of [17] under the UH condition as follows. For each patch $U_{k}^{(i)}$, there exist a $C_{U_{k}^{(i)}}>0$ such that
$\max_{q_{1},\dots,q_{w}\in\{1,\dots,N\}}\sup_{y\in U_{k}^{(i)}}\left|\frac{\partial}{\partial x_{k_{1}}}\cdots\frac{\partial}{\partial x_{k_{w}}}P^{X}_{t}f(y)\right|\\ \leq C_{U_{k}^{(i)}}t^{-{(w-1)l}/{2}}\sup_{y\in U_{k}^{(i)}}\left|\nabla f(y)\right|_{\infty}.$ (32)
Then, using the Taylor expansion, there exists a polynomial $q_{U_{k}^{(i)}}$ of order $r+1$ such that
$\sup_{y\in U_{k}^{(i)}}\left|P^{X}_{t}f(y)-q_{U_{k}^{(i)}}(y)\right|\\ \leq\operatorname{rad}\left(U_{k}^{(i)}\right)^{r+1}\sup_{y\in U_{k}^{(i)}}\max_{q_{1}+\dots+q_{N}=r+1}\left|\frac{\partial^{q_{1}}}{\partial x_{1}^{q_{1}}}\cdots\frac{\partial^{q_{N}}}{\partial x_{N}^{q_{N}}}f(y)\right|$ (33)
for $y_{0}\in U_{k}^{(i)}$. By using the property of reduced measure, we have
$\begin{split}&\sum_{i=1}^{n-1}\sup_{x_{0}\in\mathbb{R}^{N}}\left|\nu^{(i)}\left(P^{X}_{T-t_{i}}f\right)-\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\left(P^{X}_{T-t_{i}}f\right)\right|\\ &\leq\sum_{i=1}^{n-1}\sup_{x_{0}\in\mathbb{R}^{N}}\sum_{j=1}^{L_{i}}\operatorname{rad}\left(U_{k}^{(i)}\right)^{m+1}\left(\max_{k_{1}+\dots+k_{N}=m+1}\left.\nu^{(i)}\right|_{U_{j}^{(i)}}\left(\frac{\partial^{k_{1}}}{\partial x_{1}^{k_{1}}}\cdots\frac{\partial^{k_{N}}}{\partial x_{N}^{k_{N}}}P^{X}_{T-t_{i}}f\right)\right.\\ &\qquad\qquad\qquad\left.-\max_{k_{1}+\dots+k_{N}=m+1}\left.\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\right|_{U_{j}^{(i)}}\left(\frac{\partial^{k_{1}}}{\partial x_{1}^{k_{1}}}\cdots\frac{\partial^{k_{N}}}{\partial x_{N}^{k_{N}}}P^{X}_{T-t_{i}}f\right)\right)\\ &\leq\sum_{i=1}^{n-1}\sum_{j=1}^{L_{i}}\operatorname{rad}\left(U_{k}^{(i)}\right)^{m+1}\operatorname{wt}\left(U_{k}^{(i)}\right)\max_{k_{1}+\dots+k_{N}=m+1}\sup_{y\in U_{k}^{(i)}}\left|\frac{\partial^{k_{1}}}{\partial x_{1}^{k_{1}}}\cdots\frac{\partial^{k_{N}}}{\partial x_{N}^{k_{N}}}P^{X}_{T-t_{i}}f(y)\right|\end{split}$
Using (32), we have
$\begin{split}&\sum_{i=1}^{n-1}\sup_{x_{0}\in\mathbb{R}^{N}}\left|\nu^{(i)}\left(P^{X}_{T-t_{i}}f\right)-\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\left(P^{X}_{T-t_{i}}f\right)\right|\\ &\qquad\leq C_{2}(T)\sum_{i=1}^{n-1}\frac{\sum_{j=1}^{L_{i}}\operatorname{rad}\left(U_{k}^{(i)}\right)^{m+1}\operatorname{wt}(U_{k}^{(i)})}{(T-t_{i})^{m/2}}\|\nabla f\|_{\infty}\end{split}$
thus. ∎
Example 2 (Refined patch conditon).
For patches $\left\{U_{k}^{(i)}\right\}_{\begin{subarray}{c}i=1,\dots,\sharp\Delta,\\ k=1,\dots,l_{i}\end{subarray}}$ that satisfy
$\operatorname{rad}\left(U_{k}^{(i)}\right)=(T-t_{i})^{{m}/(2(m+1))}\left(l_{i}\operatorname{wt}\left(U_{k}^{(i)}\right)\right)^{-{1}/{(m+1)}}s_{i}^{{1}/{2}},$ (34)
we have the recombination error of order $O(n^{-(m-1)/{2}})$.
3.2. patch division algorithm
In this subsection,we present a recursive algorithm for constructing a set of patches $\left\{U_{k}^{(i)}\right\}_{\begin{subarray}{c}i=1,\dots,\sharp\Delta,\\ k=1,\dots,l_{i}\end{subarray}}$ that satisfies (34) from the probability measure $\nu^{(i)}=\sum_{j=1}^{m_{i}}\hat{w}_{j}^{(i)}\delta_{p^{(i)}_{j}}$. To this end, given a measure $\bar{\nu}=\sum_{j=1}^{m}w_{j}\delta_{p_{j}}$ on $\mathbb{R}^{N}$, we define a penalty function $\Psi_{i}$ as
$\Psi_{i}\left(\bar{\nu}\right)=\operatorname{rad}(U^{\bar{\nu}})-(T-t_{i})^{{m}/({2(m+1)})}{s_{i}}^{{1}/{2}}\left(\operatorname{wt}(U^{\bar{\nu}})\right)^{{1}/{(m+1)}},$ (35)
where $U^{\bar{\nu}}=\operatorname{supp}(\bar{\nu})=\left\{p_{j}\right\}_{j=1}^{m}$. The penalty function corresponds to the difference between the radius and the patch condition given in Example 2.
Starting from $\nu^{(i)}=\sum_{j=1}^{m_{i}}\hat{w}_{j}^{(i)}\delta_{p_{j}^{(i)}}$, we recursively execute the following steps 1–3 to construct a collection of patches based on the penalty function. More precisely, given a measure $\bar{\nu}$, we perform Steps 1–3 to divide it into two measures $\bar{\nu}_{1}$ and $\bar{\nu}_{2}$, and continue the procedure recursively for each measure until the penalty function $\Psi_{i}(\bar{\nu}_{j})$ becomes positive:
$\Psi_{i}(\bar{\nu}_{j})>0\quad(j=1,2).$
If the penalty function becomes positive for a measure $\bar{\nu}_{j}$, we define the support of $\bar{\nu}_{j}$ as a patch $U_{k}^{(i)}$, and increment $k$ by one.
  1. Step 1.
    Calculate the projections of the nodes onto the least-squares direction We set the starting point $c$ of the projected vector as the centre of mass of the given nodes. Thus we first calculate the centre of mass $c=\begin{pmatrix}c^{(1)}&\cdots&c^{(N)}\end{pmatrix}\in\mathbb{R}^{N}$, that is
    $c=\sum_{j=1}^{m}w_{j}p_{j}.$ (36)
    Then, we set the end point of the projected vector $e\in\mathbb{R}^{N}$ as the minimiser of weighted squared error, that is
    $e=\operatorname*{arg\,min}_{c\in\mathbb{R}^{N}}\sum_{j=1}^{m}w_{j}\left\|p_{j}-c\right\|_{\mathbb{R}^{N}}.$ (37)
  2. Step 2.
    Sort the nodes $\left\{\left(p_{j},w_{j}\right)\right\}_{j=1,\dots,m}$ in order of inner product with the projected vector, that is
    $\left\langle p_{j}-c,e-c\right\rangle$ (38)
    to $\left\{\left(p_{s(j)},w_{s(j)}\right)\right\}_{j=1,\dots,m}$.
  3. Step 3.
    Divide the nodes into two patches so that each patch has an equal number of nodes. Namely, divide $\bar{\nu}$ into two measures $\bar{\nu}_{1}$ and $\bar{\nu}_{2}$ as
    $\bar{\nu}_{1}=\sum_{j=1}^{\lfloor m/2\rfloor}w_{s(j)}\delta_{p_{s(j)}},\;\;\text{and}\;\;\bar{\nu}_{2}=\sum_{j=\lfloor m/2\rfloor+1}^{m}w_{s(j)}\delta_{p_{s(j)}}.$ (39)
Figures 2 and 3 show an example of patch division. This example is in the case of pricing an Asian option under the Heston model, which we discuss in subsection 4.1, and the number of partitions $n=6$ and the fifth times step $t_{5}=(5/6)T$. Here, the sizes represent the weights of the nodes, the colors represent which patch the node is included in. We can see that nodes with smaller weights, that is smaller size nodes, are divided to one large patch.
Remark 4 (Advantages of the above-described patch division algorithm).
The above algorithm enables a patch division that satisfies the condition in (34), while minimizing the total number of patches as much as possible. Moreover, due to its recursive nature, the patch division can be executed efficiently. This reduces the overall computational cost of the recombination method.
Figure 2. Example of patch division
Figure 3. Example of patch division (from a different angle)
4. Numerical results
To discuss the practical efficiency of the high-order recombination, we conduct the following numerical experiments.
4.1. Convergence of Discretisation error
First, to verify the theoretical converges in practical problems, we consider the pricing an Asian call option under the Heston model, which is also considered in [24, 20, 26]. Let $Y^{(1)}(t,x)$ be the price process and $Y^{(2)}(t,x)$ be the volatility process as follows.
$\begin{split}&dY^{(1)}(t,x)=\mu Y^{(1)}(t,x)\;\mathrm{d}t+Y^{(1)}(t,x)\sqrt{Y^{(2)}(t,x)}\;\mathrm{d}B^{(1)}(t),\\ &dY^{(2)}(t,x)=\alpha(\theta-Y^{(2)}(t,x))\;\mathrm{d}t+\beta\sqrt{Y^{(2)}(t,x)}\left(\rho\;\mathrm{d}B^{(1)}(t)+\sqrt{1-\rho^{2}}\;\mathrm{d}B^{(2)}(t)\right)\end{split}$
where $x=(x_{1},x_{2})\in(\mathbb{R}_{>0})^{2}$, $-1\leq\rho\leq 1$. We set the parameters $\alpha,\beta,\theta$ to satisfy the Feller condition:
$02\alpha\theta-\beta^{2}>$ 0 (40)
to ensure the existence and uniqueness of a solution to the CIR process $Y^{(2)}(t,x)$ [5]. Note that when the CIR process is discretised using the N–V method, the positivity can be guaranteed under the following condition:
$4\alpha\theta-\beta^{2}>0,$ (41)
which is weaker than (40). This condition for the positivity can be derived the discretisation formula presented in [24]. The payoff of the Asian call option on this asset with maturity $T$ and strike $K$ is $\max(Y^{(3)}(T,x)/T-K,0)$ where
[…]
Let $Y(t,x)={}^{t}(Y^{(1)}(t,x),Y^{(2)}(t,x),Y^{(3)}(t,x))$. Here we change the coordinate of $Y^{(2)}$ for numerical stability of the simulation. Then $Y(t,x)$ is the solution of the Stratonovich form SDEs
$Y(t,x)=\sum_{i=0}^{2}\int_{0}^{t}W_{i}\left(Y(s,x)\right)\circ\mathrm{d}B^{i}(s)$
where
$W_{0}\left(\begin{array}[]{c}y_{1}\\ y_{2}\\ y_{3}\end{array}\right)$ $=\left(\begin{array}[]{c}y_{1}\left(\mu-{y_{2}}/{2}\right)-{\beta\rho}y_{1}/4\\ \alpha\left(\theta-y_{2}\right)-{\beta^{2}}/{4}\\ y_{1}\end{array}\right),$
$W_{1}\left(\begin{array}[]{c}y_{1}\\ y_{2}\\ y_{3}\end{array}\right)$ $=\left(\begin{array}[]{c}y_{1}\sqrt{y_{2}}\\ \beta\rho\sqrt{y_{2}}\\ 0\end{array}\right),$
$W_{2}\left(\begin{array}[]{c}y_{1}\\ y_{2}\\ y_{3}\end{array}\right)$ $=\left(\begin{array}[]{c}0\\ \beta\sqrt{y_{2}(1-\rho^{2})}\\ 0\end{array}\right).$
We set $T=1,K=1.05,\mu=0.05,\alpha=2.0,\beta=0.1,\theta=0.09,\rho=0.3$, and $(x_{1},x_{2})=(1.0,0.09)$ and take the true value as follows:
[…]
This value is obtained by the third order method [26] and the quasi-Monte Carlo with the number of partitions $n=100$ and the number of sample points $M=10^{9}$. For more verification, refer to Section 4 of [26].
Figure 4 shows the convergences of the total error
$\left|E_{\mu_{x,\Delta,\mathcal{U}}^{(n)}}\left[f\left(Y^{(n)}(T,x)\right)\right]-E_{P}\left[f\left(Y(T,x)\right)\right]\right|,$
which is the difference between the expectations under the recombined measure $\mu_{x,\Delta,\mathcal{U}}^{(n)}$ and the original measure $P$, with respect to the numbers of partition $n$. Here, we set a time partition $\Delta$ as the even partition, that is $s_{i}={1}/{n}$, and a set of radii as in Corollary 2. See Remark 6 below on the reasons for adopting the even partition here.
The errors in the E–M and N–V method with and without the recombination are plotted. Figure 5 shows the growth of the number of support. As clearly shown in Figure 4, the theoretical convergences of error can be attained in each method, whether we use the recombination or not. If we do not use the recombination, that is the whole tree calculation, only executable up to around $n=8$, these are limits due to the computational complexity. On the other hand, if we use the recombination, we can calculate more larger numbers of partitions $n$, and we obtain the theoretical order. In fact, Figure 5 clearly shows that the recombination enables us to suppress the exponential growth of the number of support to the polynomial growth, $O(n)$ to $O(n^{2})$ in our example. This indicates that the proposed recombination algorithm enables us to conquer the problem of the escalating computational complexity in this practical problem.
Remark 5 (Heston model and the UH condition).
The Heston model do not satisfy the UH condition which we assume in Theorem 5. Nevertheless, the numerical result demonstrates that our algorithm works for the Heston model case. In fact, the following considerations generally justify the numerical result. In the Heston model, the positivity of the discretised process by the N–V method ensure that $\operatorname{supp}(\nu^{(i)})\subset\mathbb{R}^{3}_{>0}=\left\{(y_{1},y_{2},y_{3})\;|\;y_{1},y_{2},y_{3}>0\right\}$, where $\nu^{(i)}$ is an intermediate measure defined in (11). This indicates that for any patch division $\left\{U_{k}^{(i)}\right\}_{k=1,\dots,l_{i}}$ of $\operatorname{supp}(\nu^{(i)})$ there exists an $\varepsilon>0$ such that $\bigcup_{k=1,\dots,l_{i}}U_{k}^{(i)}\subset\left\{(y_{1},y_{2},y_{3})\;|\;y_{1},y_{2},y_{3}>\varepsilon\right\}$. Moreover, in the Heston model, $\left\{W_{1},W_{2},[W_{0},W_{1}]\right\}$ spans >$\mathbb{R}^{3}_{>0}$, thus, Theorem 2 of [12] implies that for each patch $U_{k}^{(i)}$ there exists a $C_{U_{k}^{(i)}}$ such that (32) holds, and
$\max_{\begin{subarray}{c}i=1,\dots,n\\ k=1,\dots,l_{i}\end{subarray}}C_{U_{k}^{(i)}}<\infty.$
However, this $\varepsilon$ depends on the discretisation and recombination algorithm, so we cannot know the value of $\varepsilon$ before executing the algorithm. Therefore, to fully justify our algorithm for the Heston model case, we need to prove Theorem 5 under a weaker condition than the UH condition such as the UFG condition, which is a subject for future work.
Remark 6 (Optimality of even partitioning).
Throughout these numerical experiments, we use the even partition, that is $s_{i}={T}/{n}$, which is proved to be optimal for the original continuous versions of N–V and N–N methods in [14] and not for the discrete versions. However, our numerical experiments show that it also works for the discretised N–V method, as similarly demonstrated in [22].
Figure 4. Total Error (Heston)
Figure 5. # supp(Heston)
4.2. Impact of patch division algorithm on recombination error
To investigate the effect of patch division on recombination error, we conduct further validation. We introduce a scaling parameter $\kappa$ to the penalty function as follows.
$\Psi^{(\kappa)}(U)=\text{rad}(U)-\kappa(T-t_{i})^{{m}/({2(m+1)})}s_{i}^{{1}/{2}}\left(\operatorname{wt}(U)\right)^{{1}/(m+1)}.$ (42)
For a fixed the number of partition $n$, by varying $\lambda$, we calculate the recombination errors
$\left|E_{\mu_{x,\Delta,\mathcal{U},\kappa}^{(n)}}\left[f\left(Y^{(n)}(T,x)\right)\right]-E_{\lambda_{x,\Delta,\mathcal{U}}^{(n)}}\left[f\left(Y(T,x)\right)\right]\right|,$
namely the difference between the expectations under the recombined measure $\mu_{x,\Delta,\mathcal{U},\kappa}^{(n)}$ constructed by using the penalty function (42) and the measure without recombining $\lambda_{x,\Delta,\mathcal{U}}^{(n)}$, that is the whole tree calculation defined by (9), for different number of support. We compare our patch division algorithm (Recursive_Refcond) with the concentric circle dividing (CC), the random dividing (Rand), and the recursive path dividing with Litterer–Lyons condition (Recursive_LLcond). Note that the theoretical patch division condition is not necessarily satisfied in CC and Rand.
Figure 6 shows the recombination errors in pricing the Asian option in the Heston model, as the same setting as in the previous subsection. Here, we fix the number of partitions $n=8$. It seems that convergence efficiencies of each patch division are as follows:
${\rm Recursive\_Refcond}>{\rm Recursive\_LLcond}>>{\rm Rand}>{\rm CC}.$
Figure 6. Total Error (Heston)
In addition, to investigate in which cases the patch conditions have a greater influence on the recombination error, we conduct a validation on the spiky payoff cases. Setting $\left(\tilde{X}(T,x),\tilde{Y}(T)\right)=\left(Y_{3}(T,x)/Y_{1}(0,x),Y_{2}(T,x)/Y_{2}(0,x)\right)$ and
$f\left(\tilde{x},\tilde{y}\right)=h-\frac{h}{r}\sqrt{(x-x_{c})^{2}+(y-y_{c})^{2}},$
where $h$ is the height, $r$ is the radius, and $(x_{c},y_{c})$ is the centre of cone. We set the height as $h=\frac{10}{r^{2}}$so that the volume of the cones remain constant. In this validation, We consider the following 80 cone payoffs in this coordinate.
  • Radius : 5.0, 4.0, 3.0, 2.0, 1.5, 1.2, 1.0, 0.9
  • Centre of cone : (0,0), (1,0), (0,1), (1,1), (2,0), (0,2), (2,2), (3,0), (0,3), (3,3)
Figure 7 shows some of above payoffs and nodes at the maturity, which indicate the above settings are exhaustive. Fixing the number of partitions $n=6,8$, we compare the recombination errors with respect to the number of supports at the maturity in each patch division method.
Figures 8 and 9 summarize the averages of slopes
$\frac{\left|E_{\mu_{x,\Delta,\mathcal{U},\kappa}^{(n)}}\left[f\left(Y(^{(n)}T,x)\right)\right]-E_{\lambda_{x,\Delta,\mathcal{U}}^{(n)}}\left[f\left(Y(T,x)\right)\right]\right|}{\operatorname{card}\left(\operatorname{supp}\left(\mu_{x,\Delta,\mathcal{U},\kappa}^{(n)}\right)\right)}$
with its confidence intervals. These graphs indicate that our patch division algorithm works well for most cases.
Figure 7. Spiky payoff
Figure 8. Average of slopes and its confidence interval
Figure 9. Average of slopes and its confidence interval
Appendix1: Construction of a reduced measure
Let $\nu$ be an induced probability measure on a patch $U$, defined in (14). $\mathcal{N}_{\nu}$ denotes the number of nodes included in the patch $U$. Let $G=\{g_{1},g_{2},\dots,g_{\mathcal{M}_{G}}\}$ be a set of test functions, where $\mathcal{M}_{G}$ is the number of test functions and each $g_{i}:\tilde{\Omega}\rightarrow\mathbb{R}$ is a measurable function. Hereinafter, $x^{(j)}$ denotes the $j$th coordinate of $x$, that is $x=\left(x^{(1)},\dots,x^{(N)}\right)$.
  1. (1)
    Calculate the basis of kernel of the test function matrix.
    We define $\mathcal{M}_{G}\times\mathcal{N}_{\nu}$ matrix $A_{0}$ as follows:
    $A_{0}=\left(\begin{array}[]{cccc}g_{1}\left(y_{1}\right)&g_{1}\left(y_{2}\right)&\ldots&g_{1}\left(y_{\mathcal{N}_{\nu}}\right)\\ g_{2}\left(y_{1}\right)&&\ldots&\\ \vdots&\vdots&\ddots&\vdots\\ g_{\mathcal{M}_{G}}\left(y_{1}\right)&&\ldots&g_{\mathcal{M}_{G}}\left(y_{\mathcal{N}_{\nu}}\right)\end{array}\right)$ (43)
    Then, we calculate ${\rm ker}(A_{0})$ using the singular value decomposition. Since ${\rm dim}({\rm ker}(A_{0}))=\mathcal{N}_{\nu}-\mathcal{M}_{G}^{\prime}$ where $\mathcal{M}_{G}^{\prime}={\rm rank}(A_{0})$, we can choose a basis for ${\rm ker}(A_{0})$ consisting of $\mathcal{N}_{\nu}-\mathcal{M}_{G}^{\prime}$ elements denoted as $\left\{v_{0,1},v_{0,2}\dots,v_{0,\mathcal{N}_{\nu}-\mathcal{M}_{G}^{\prime}}\right\}$. We also let $\beta_{0}=\prescript{t}{}{\begin{pmatrix}\phi_{1}&\cdots&\phi_{\mathcal{N}_{\nu}}\end{pmatrix}}\in[0,1]^{\mathcal{N}_{\nu}}$.
  2. (2)
    Eliminate the nodes to be recombined and update the weights.
    Determine the set of nodes {}$\left\{\right\}_{i=1}^{\mathcal{N}_{\nu}-\mathcal{M}^{\prime}_{G}}$ that are to be eliminated by the procedures specified for each $i=1,\dots,\mathcal{N}_{\nu}-\mathcal{M}^{\prime}_{G}$ as follows:
    • (2-1)
      Determine a node to be eliminated by specifying the number of all nodes to be eliminated.
      $\alpha_{i}=\min_{1\leq j\leq\mathcal{N}_{\nu}-i+1}\left\{\left.\frac{\beta_{i-1}^{(j)}}{\left(v_{(i-1),1}\right)^{(j)}}\;\right|\;\left(v_{(i-1),1}\right)^{(j)}>0\right\}$ (44)
      We also let
      $e_{i}=\operatorname*{arg\,min}_{1\leq j\leq\mathcal{N}_{\nu}-i+1}\left\{\left.\frac{\beta_{i-1}^{(j)}}{\left(v_{(i-1),1}\right)^{(j)}}\;\right|\;\left(v_{(i-1),1}\right)^{(j)}>0\right\},$ (45)
      that is $\alpha_{i}=\displaystyle{\frac{\beta_{i-1}^{(e_{i})}}{\left(v_{(i-1),1}\right)^{(e_{i})}}}$.
    • (2-2)
      Eliminate the node $(y_{e_{i}},\phi_{e_{i}})$.
      Eliminating the $e_{i}$th column of $A_{i-1}$, we obtain $A_{i}$ as an $\mathcal{M}_{G}\times(\mathcal{N}_{\nu}-i)$ matrix. Then, we update the probability measure by defining $\beta_{i}\in[0,1]^{\mathcal{N}_{\nu}-i}$ as follows:
      $\beta_{i}^{(j)}=\left\{\begin{array}[]{c}\beta_{i-1}^{(j)}-\alpha_{i}\left(v_{(i-1),1}\right)^{(j)}\quad\text{for $j=1,\dots,e_{i}-1$},\\ \beta_{i-1}^{(j+1)}-\alpha_{i}\left(v_{(i-1),1}\right)^{(j+1)}\quad\text{for $j=e_{i},\dots,\mathcal{N}_{\nu}-i$.}\end{array}\right.$ (46)
    • (2-3)
      Update the basis of kernel of the test function matrix.
      For $l=1,\dots,\mathcal{N}_{\nu}-\mathcal{M}_{G}^{\prime}-i$, we define
      $\begin{split}&d_{i,l+1}=\frac{\left(v_{(i-1),(l+1)}\right)^{(e_{i})}}{\left(v_{(i-1),1}\right)^{(e_{i})}},\\ &v_{i,l}=v_{(i-1),(l+1)}-d_{i,l+1}v_{(i-1),1}\in\mathbb{R}^{\mathcal{N}_{\nu}},\\ &\left(v_{i,l}\right)^{(j)}=\left\{\begin{array}[]{c}\left(v_{i,l}\right)^{(j)}\;\quad\text{for $j=1,\dots,e_{i}-1$}\\ \left(v_{i,l}\right)^{(j+1)}\;\quad\text{for $j=e_{i},\dots,\mathcal{N}_{\nu}-i$}\end{array}\right.\end{split}$ (47)
Example 3 (An Illutrative example).
Let us give a simple example. Let $X,Y$ be $2$-dimensional discrete random variables on $\Omega=\{\omega_{0},\dots,\omega_{7}\}$ such that
$\begin{split}&(X,Y)(\omega_{0})=(0,0),\;\;(X,Y)(\omega_{1})=(1,0),\;\;(X,Y)(\omega_{2})=(0,1),\;\;(X,Y)(\omega_{3})=(1,1),\\ &(X,Y)(\omega_{4})=(2,0),\;\;(X,Y)(\omega_{5})=(2,1),\;\;(X,Y)(\omega_{6})=(2,2),\;\;(X,Y)(\omega_{7})=(1,2)\end{split}$
and $\tilde{Q}$ be a probability measure on $\Omega$ such that $\tilde{Q}(\omega_{0})=\dots=\tilde{Q}(\omega_{7})=1/8$. From $\tilde{Q}$, find its reduced measure $Q$ such that the moments of this $(X,Y)$ is invariant up to the second order, that is to say $\tilde{Q}$ is the reduced measure from $Q$ with respect to $G=\left\{g_{i}(X,Y)\right\}_{i=1}^{6}$ where $\left\{g_{i}\right\}_{i=1}^{6}\subset\mathbb{R}[x,y]$ is defined as
$g_{1}(x,y)=x^{0}y^{0}=1,\quad g_{2}(x,y)=x,\quad g_{3}(x,y)=y,$
$g_{4}(x,y)=x^{2},\quad g_{5}(x,y)=y^{2},\quad g_{6}(x,y)=xy.$
The problem reduces to finding a basis of $\operatorname{Ker}{A}$ of the following matrix $A$, where the $i$th column and the $j$th row corrrespond to the state $\omega_{i-1}$ and the monomial $g_{j}$ respectively:
$A=\begin{pmatrix}1&1&1&1&1&1&1&1\\ 0&1&0&1&2&2&2&1\\ 0&0&1&1&0&1&2&2\\ 0&1&0&1&4&4&4&1\\ 0&0&1&1&0&1&4&4\\ 0&0&0&1&0&2&4&2\end{pmatrix}.$
Then we see that $\operatorname{Ker}(A)=\mathbb{R}v_{1}+\mathbb{R}v_{2}$ where
$\begin{split}v_{1}&={\prescript{t}{}{\begin{pmatrix}-1&2&1&-2&-1&1&0&0\end{pmatrix}}},\\ v_{2}&={\prescript{t}{}{\begin{pmatrix}2&-3&-2&2&1&0&-1&1\end{pmatrix}}}.\end{split}$
If we set
$\begin{split}Q&=\tilde{Q}-(1/8)v_{1}-(1/24)v_{2}\\ &=\prescript{t}{}{\begin{pmatrix}{1}/{6}&{0}&{1}/{12}&{7}/{24}&{5}/{24}&{0}&{1}/{6}&{1}/{12}\end{pmatrix}},\end{split}$
we see that $Q$ satisfies
$\begin{split}&E^{\tilde{Q}}[X]=E^{Q}[X],\;\;E^{\tilde{Q}}[Y]=E^{Q}[Y],\;\;\\ &E^{\tilde{Q}}[X^{2}]=E^{Q}[X^{2}],\;\;E^{\tilde{Q}}[XY]=E^{Q}[XY],\;\;E^{\tilde{Q}}[Y^{2}]=E^{Q}[Y^{2}].\end{split}$
Appendix2: Pseudo code of the high-order recombination algorithm
Let $0=t_{0}<t_{1}<\dots<t_{n}=T$ be a partition of time $[0,T]$.
1
2
Input : Number of partitions : $n$,
Parameters of SDE : $x_{0}$, $V_{0},\dots,V_{d}$
Option parameters : maturity $T$, payoff function $f$
Output : $E_{\mu_{t_{n}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}}f(X(T,x_{0}))$
3 Struct Node contains
       $values$ ;
        /* Realized value $\hat{X}(t_{i},x_{0}))(\omega_{j}\in\Omega_{i}$) */
       $weight$ ;
        /* Weight (a value of measure at this node $P\left(\omega_{j}\right)$) */
       $patchnum$ ;
        /* patch number of this node */
4
5 end
6 Function Main():
7       $leaves\leftarrow\text{array of {\it Node}}$;
8       $nextleaves\leftarrow\text{array of {\it Node}}$;
9
      $leaves[0].value\leftarrow x_{0}$ ;
        /* Set the initial value */
10       $leaves[0].weight\leftarrow 1.0$;
11
12      for $i\leftarrow 1$ to $n$ do
13             $nextleaves\leftarrow\textnormal{{OneStepFwd}}(leaves,i,n,T,V_{0},\dots,V_{d})$;
14             $numpatch\leftarrow 0$;
15             PatchDiv $(nextleaves,numpatch)$;
16             $nextleaves\leftarrow$ Recombination $(nextleaves)$;
17             $leaves\leftarrow nextleaves$;
18
19       end for
20      $nextleaves\leftarrow\textnormal{{OneStepFwd}}(leaves,i,n,T,V_{0},\dots,V_{d})$;
21       $value\leftarrow\text{Payoff}(nextleavesleaves,f)$;
22       return value;
23
24
1 Function OneStepFwd($leaves,i,n,T,V_{0},\dots,V_{d}$):
        /* Calculate $\hat{X}^{(t_{i+1},x_{0})}$ from $\hat{X}^{(t_{i},x_{0})}$, using (23) */
2       $nextleaves\leftarrow\text{array of {Node}}$;
3       forall $node$ in $leaves$ do
              /* Generate RVs $\{\eta_{i}\}$ and solve ODEs */
4             for $k\leftarrow 1$ to $2\times 3^{d}$ do
5                   nextval $\leftarrow$ Solve ODEs along with $V_{0},\dots,V_{d}$;
                    /* Using (23) */
6                   nextleaves.node $\leftarrow$ nextval;
7
8             end for
9
10       end forall
11      return $nextleaves$
12
[…]
  /* Equation   */
1;
2 if $Cond==True$ then
3       forall $node$ in leaves do
4             $leaves.node.patchnum\leftarrow numpatch$;
5
6       end forall
7      $*numpatch\leftarrow*numpatch+1$;
8
9 end if
10else
11       SortNodes ($leaves$);
12       $leftleaves\leftarrow leaves[numnode/2]$;
13       $rightleaves\leftarrow leaves[numnode/2:numnode]$;
14       PatchDiv ($leftleaves$, numpatch);
15       PatchDiv ($rightleaves$, numpatch);
16
17 end if
18
19 Function SortNodes($leaves$):
20       $g\leftarrow\text{gravitypoint}(leaves)$;
        /* Using (36) */
21       $e\leftarrow\text{endpoint}(leaves)$;
        /* Using (37) */
22       forall $node$ in $leaves$ do
23             $node.order\leftarrow\text{innerprod}(node.value,g,e)$;
              /* Using (38) */
24
25       end forall
26      QuickSort($leaves$, $leaves.node.order$);
        /* Using external library */
27
28
1 Function SortNodes($leaves$):
2       $g\leftarrow\text{gravitypoint}(leaves)$;
        /* Using (36) */
3       $e\leftarrow\text{endpoint}(leaves)$;
        /* Using (37) */
4       forall $node$ in $leaves$ do
5             $node.order\leftarrow\text{innerprod}(node.value,g,e)$;
              /* Using (38) */
6
7       end forall
8      QuickSort($leaves$, $leaves.node.order$);
        /* Using external library */
9
10